# Fluctuations of Young diagrams for symplectic groups and semiclassical orthogonal polynomials

### v2

@@ 3.1. Double-scaling asymptotic for Krawtchouk polynomials
$n=HK,\quad m=gK+s\sqrt{K}+l,\quad K\to\infty,$
where $g,s,l$ are fixed parameters, $10<g<$ 1, and the critical point is located at $g=\frac{1}{2}$.
Numerous studies have explored the asymptotic behavior of the Krawtchouk polynomials $\widetilde{K}_{m}(i/n)$ using methods such as steepest descent and the Riemann-Hilbert approach (see [6], [8], [9]). In our analysis, we aim to extend these computations to the regime where the lattice variable $i$ remains unbounded in the double-scaling limit.
To ensure that $i$ is not constrained, we introduce a new variable $\mu$ defined as
$i=\frac{K}{2}(1-2\mu),\quad i\in[-K/2,K/2],$
where $10<\mu<$ 1 and $\mu\neq\frac{1}{2}$. Now let us consider on the asymptotic expansion of
$\widetilde{K}_{m}(i/n)=(-1)^{m}\frac{m!}{n^{m}}\int_{\Gamma}(1-(1-p)z)^{K/2-i}(1+pz)^{K/2+i}z^{-m}\frac{\mathrm{dz}}{2\pi iz},$ (3.1)
where the contour $\Gamma$ is a positive oriented closed unity loop that encloses 0, $|z|=1$. Changing the variable $t=1/(1-(1-p)z)$ one can find the integral
$I=\int_{\Gamma}(1-(1-p)z)^{K/2-i}(1+pz)^{K/2+i}z^{-m}\frac{\mathrm{dz}}{2\pi iz}=$
$=-(-1)^{m+1}\frac{(1-p)^{m-i-K/2}}{2\pi i}\int_{\Gamma^{\prime}}\frac{t^{s\sqrt{K}+l-1}dt}{(1-t)^{s\sqrt{K}+l+1}}\exp\left\{K\ln{\frac{(t-p)^{1-\mu}}{(1-t)^{g}}}t^{g-1}\right\},$
where the contour $\Gamma^{\prime}$ is a smooth negative -oriented circle around 1. Now we can use saddle-point method to analyze the asymptotic behavior of the integral (see [43])
[…]
$\theta_{\text{SDP}}^{j}=-\frac{\alpha_{j}}{2}\pm\frac{\pi}{2},\quad\alpha_{j}=\arg(f^{\prime\prime}(t_{0}^{j})),$
with
$G(t)=\frac{t^{s\sqrt{K}+l-1}}{(1-t)^{s\sqrt{K}+l+1}},$
$f(t)=(1-\mu)\ln(t-p)+(g-1)\ln t-g\ln(1-t),$
where the $\theta_{\text{SDP}}^{j}=\arg\left(\frac{dt}{d\rho}\right)_{\rho=0}$ is the tangent angle of the steepest descent path in the saddle point $t_{0}^{j}$, $\rho$ is the parameter along the minimax contour and the sign in the expression of $\theta_{\text{SDP}}^{j}$ defined by the steepest descent path orientation in the saddle points.
@@
So exactly up to factor $\frac{1}{2\mu}$ and changing $\mu\to p$, roots of the action in the considered limit the same as the roots of the generalized Krawtchouk polynomials with $p\neq 1/2$, described above. Starting from the region of two complex-conjugate roots $1/2-\sqrt{\mu(1-\mu)}<g<1/2+\sqrt{\mu(1-\mu)}$ for a fixed $\mu$ (for a fixed $g$, this region is defined by the inequality $1/2-\sqrt{g(1-g)}<\mu<1/2+\sqrt{g(1-g)}$)  [^4] This is also a direct consequence of the self-duality of the Krawtchouk polynomials., let us choose the parametrization $g=1/2+d\sqrt{\mu(1-\mu)},\;d\in[-1,1]$. Then one can easily define two branches of an algebraic rational plane curve $t_{0}(d;\mu)=x(d;\mu)+iy_{\pm}(d;\mu)$
@@
Let us focus on the case of two complex conjugate roots. One can express all main values from the equation (3.2)
@@
$G(t_{0}^{\pm})=\frac{1}{(t_{0}^{\pm})^{2}}\left(\frac{t_{0}^{\pm}}{1-t_{0}^{\pm}}\right)^{s\sqrt{K}+l+1}=\frac{2\mu}{g}\left(\sqrt{\frac{1-g}{g}}\right)^{s\sqrt{K}+l-1}e^{\pm i\big{[}\zeta_{2}(s\sqrt{K}+l-1)-2\tau\big{]}},$ (3.6)
@@
$\quad+\pi\cdot\Theta\left(\frac{1}{2}-\frac{\sqrt{\mu}}{2}-g\right)-\pi\cdot\Theta\left(g-\left(\frac{1}{2}+\frac{\sqrt{\mu}}{2}\right)\right)\bigg{]}.$ (3.7)
@@
Functions $\zeta_{1}$ and $\zeta_{2}$ are the angles of these shifts relative to $\chi=\pm\arg(t_{0}^{\pm}(1/2))$
@@
The $\sigma$ is defined from the denominator of $-\pi\leq\arg f^{\prime\prime}(t_{0}^{\pm})\leq\pi$ as
@@
for $t\in\mathbb{C},\,\mu,g\in\mathbb{R},\,\mu\in(0,1),\,g\in(1/2-\sqrt{\mu(1-\mu)},1/2+\sqrt{\mu(1-\mu})$, and
@@
The solutions of these equations are symmetric relative to the reflection $t\to\overline{t}$, it follows from the logarithmicity of the action $f(t)$. By the symmetry of the SDP branches
@@
[…]
$\sim(-1)^{m}(1/2)^{m-K/2}\frac{(\sqrt{1-g})^{m-1/2}}{(\sqrt{g})^{m+1/2}}\left(\sqrt{\frac{\mu}{1-\mu}}\right)^{K/2-j}\left(\sqrt{\frac{1-\mu}{1-g}}\right)^{K}\sqrt{\frac{2}{\pi K}}\left(\frac{(1-\mu)\mu}{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}\right)^{1/4}\cdot$
$\cdot\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m)),$
@@
[…]
$\sim\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\sqrt{\frac{2}{\pi K}}\frac{(-1)^{m}(1/2)^{m-K/2}}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}}\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m)).$
So for modified Krawtchouk polynomial unbounded double-scaling asymptotic we finally obtain
$\widetilde{K}_{m}(j/n)\bigg{|}_{n=HK,m=gK+s\sqrt{K}+l,K\to\infty}\sim$
$\sim\frac{m!}{n^{m}}\sqrt{\frac{2}{\pi K}}\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\frac{(1/2)^{m-K/2}\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m))}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}},$ (3.11)
where $\theta_{\text{SDP}}^{-}$ and $\hat{\delta}$ are defined above as the functions of $\mu,g$.
3.2. Symplectic polynomials asymptotic
Using the equation (2.9) we can find the asymptotic behavior for $m=gK+s\sqrt{K}+l$,
$n=HK,\;K\sim\infty$ for even $m$
@@ 3.2. Symplectic polynomials asymptotic
$\cdot\bigg{(}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;m+2)\big{)}+\frac{(K-m)}{m+2}\frac{g}{1-g}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;m)\big{)}\bigg{)}+O(1/K).$ (3.12)
4. Limiting behavior of the correlation kernel
The main goal of this section is to compute the asymptotic of the correlation kernel (1.13) in following regime of the symplectic polynomials
@@ 4. Limiting behavior of the correlation kernel
with $K\to\infty$ and fixed $g,H$, and $m$ is the number of polynomial. The equation (3.12) for $m=2n$ has the following form
$G_{2n}(i/n)=g_{2n}(i/n)\sqrt{\Lambda_{2n}}=$
$=\frac{1}{(i/n)^{2}}\frac{(2n+2)!}{n^{2n+2}}F_{2n}(i)\bigg{(}\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;2n+2)\big{)}+\sin\big{(}\theta_{SDP}^{-}-\hat{\delta}(i;2n)\big{)}\bigg{)}+O(1/K),$ (4.2)
@@
with $g=2H$ and in general $20<H<1/$ 2 with $H=1/4$ for $n=k$ (square Young tableaux). Now we substitute asymptotic of the polynomials into the kernel to demonstrate its convergence to the discrete sine-kernel
@@
$\cdot\bigg{(}\sin\big{(}\theta^{-}_{\text{SDP}}(j;2n)-\hat{\delta}(j;2n)\big{)}+\sin\big{(}\theta^{-}_{\text{SDP}}(j;2n-2)-\hat{\delta}(j;2n-2)\big{)}\bigg{)}-[ij]\bigg{)}\Bigg{)}+O(1/K),$ (4.3)
@@
we aim to derive the asymptotic of $\lambda$ in the bulk regime for $-1/2<x<1/2$
@@
$\gamma,\hat{\zeta}_{1},\hat{\zeta}_{2},\hat{\tau},\arg f^{\prime\prime}(t_{0}^{-}))$. We need to choose the interval between the critical points of the action. So we can expand the piecewise functions for example in $g>1/2,\,\mu>1/2$ ($-K/2<i<0$)
@@
$\arg f^{\prime\prime}(t_{0}^{-})=\tan^{-1}\left(\frac{(2g-1)(2\mu-1)}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+2\tan^{-1}\left(\frac{-\frac{\mu}{g-1/2}+2g-1}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+\pi.$
In the other domains of parameter space such as $g>1/2,\mu<1/2$ we will get the same formulas up to ±$\pm\pi$. These contributions will cancel each other out in final answer. Next we derive the asymptotic of trigonometric part of the kernel accurate up to $O(1/K)$. It is obvious that the shift $m\to m+2$ generate the shift $g\to g+2/K$, so, by the definition of $\lambda(i;2n)$
@@
$\beta=\left(-2\zeta_{2}-\frac{\partial\hat{\zeta}_{1}}{\partial g}2+\frac{\partial\gamma}{\partial g}\right)\Bigg{|}_{m=2n,i=xK+i^{\prime}}=2\cot^{-1}\left(\frac{x}{\sqrt{2H-4H^{2}-x^{2}}}\right)+O(1/K),$
@@
$=\bigg{(}\sin\gamma_{i}+\sin\gamma_{i}\cos\beta+\cos\gamma_{i}\sin\beta\bigg{)}\bigg{(}\sin\gamma_{j}+\sin\gamma_{j}\cos\beta-\cos\gamma_{j}\sin\beta\bigg{)}-[ij]+O(1/K)=$
$=-2\sin\beta(1+\cos\beta)\sin(\gamma_{i}-\gamma_{j}),$
@@
The normalization of the kernel is not important because it can be cancelled out by the renormalization of the polynomials. Therefore we have established the convergence of the Christoffel–Darboux kernel to the discrete sine-kernel
@@
which ends the proof of Theorem 2. The limit density coincides with the result for the limit shape of symplectic diagrams obtained in [22, Thm. 5.1] after substituting $c=\lim\frac{k}{n}=\frac{1-2H}{2H}$.
@@
Figure 1. Plot of the ratio $\frac{\mathcal{K}(i,j)}{\mathcal{K}(i,i)}$ for $i=75$ obtained by sampling $500$ samples, each of $10000$ random Young diagrams inside of $n\times k$ box with $n=50,k=100$ using Proctor’s algorithm (box and whiskers plot) and its comparison to the same ratio for the Christoffel-Darboux kernel (1.9) (dashed black line) and discrete sine kernel (dotted blue) (4.4), (4.5)
5. Conclusion and future work

### v3

@@ 3 Asymptotics of the polynomials
$n=HK,\quad m=gK,\quad j\in[-K/2,K/2],\quad K\to\infty,$
where $g$ is fixed parameter, $0<g<1$, and the critical point is located at $g=\frac{1}{2}$.
Numerous studies have explored the asymptotic behavior of the Krawtchouk polynomials $\widetilde{K}_{m}(j/n)$ using methods such as steepest descent and the Riemann-Hilbert approach (see [ISMAIL1998121], [LI2000155], [Wang_Qui_2015]). In our analysis, we aim to extend these computations to the regime where the lattice variable $j$ remains unbounded in the double-scaling limit.
To ensure that $j$ is not constrained, we introduce a new variable $\mu$ defined as
$j=\frac{K}{2}(1-2\mu),\quad j\in[-K/2,K/2],$
where $0<\mu<1$ and $\mu\neq\frac{1}{2}$. Now let us consider on the asymptotic expansion of
$\widetilde{K}_{m}(j/n)=(-1)^{m}\frac{m!}{n^{m}}\int_{\Gamma}(1-(1-p)z)^{K/2-j}(1+pz)^{K/2+j}z^{-m}\frac{\mathrm{dz}}{2\pi iz},$ (3.1)
where the contour $\Gamma$ is a positive oriented closed unity loop that encloses 0, $|z|=1$. Changing the variable t=1/(1(1p)z)t=1/(1-(1-p)z) one can find the integral
$I=\int_{\Gamma}(1-(1-p)z)^{K/2-j}(1+pz)^{K/2+j}z^{-m}\frac{\mathrm{dz}}{2\pi iz}=$
$=-(-1)^{m+1}\frac{(1-p)^{m-j-K/2}}{2\pi i}\int_{\Gamma^{\prime}}\frac{dt}{t(1-t)}\exp\left\{K\ln{\frac{(t-p)^{1-\mu}}{(1-t)^{g}}}t^{g-1}\right\},$ (3.2)
where the contour $\Gamma^{\prime}$ is a smooth negative -oriented circle around 1. Now we can use steepest descent to analyze the asymptotic behavior of the integral.
Proposition 4 (see [fedoryuk1977]).
The asymptotic behavior of the contour integral of the form $\int\limits_{\Gamma}G(t)e^{Kf(t)}\mathrm{dt}$ in the limit $K\to\infty$ is given by the next series
$\int\limits_{\Gamma^{\prime}}G(t)e^{Kf(t)}dt\bigg|_{K\to\infty}=\sum\limits_{\ell}e^{Kf(t_{0}^{\ell})}\left(G(t_{0}^{\ell})\sqrt{\frac{2\pi}{K|f^{\prime\prime}(t_{0}^{\ell})|}}e^{i\theta_{\text{SDP}}^{\ell}}+O(1/K^{3/2})\right),$ (3.3)
$\theta_{\text{SDP}}^{\ell}=-\frac{\alpha_{\ell}}{2}\pm\frac{\pi}{2},\quad\alpha_{\ell}=\arg(f^{\prime\prime}(t_{0}^{\ell})),$
where the $\theta_{\text{SDP}}^{\ell}=\arg\left(\frac{dt}{d\rho}\right)_{\rho=0}$ is the tangent angle of the steepest descent path in the saddle point $t_{0}^{\ell}$, $\rho$ is the parameter along the minimax contour and the sign in the expression of $\theta_{\text{SDP}}^{\ell}$ defined by the steepest descent path orientation in the saddle points.
In our case we have
$G(t)=\frac{1}{t(1-t)},$
$f(t)=(1-\mu)\ln(t-p)+(g-1)\ln t-g\ln(1-t).$
@@
So exactly up to factor $\frac{1}{2\mu}$ and changing $\mu\to p$, roots of the action in the considered limit the same as the roots of the generalized Krawtchouk polynomials with $p\neq 1/2$, described above. Starting from the region of two complex-conjugate roots $1/2-\sqrt{\mu(1-\mu)}<g<1/2+\sqrt{\mu(1-\mu)}$ for a fixed $\mu$ (for a fixed $g$, this region is defined by the inequality $1/2-\sqrt{g(1-g)}<\mu<1/2+\sqrt{g(1-g)}$)  [^4] This is also a direct consequence of the self-duality of the Krawtchouk polynomials., let us choose the parametrization $g=1/2+d\sqrt{\mu(1-\mu)},\;d\in[-1,1]$. Then one can easily define two branches of a rational plane curve $t_{0}^{\pm}(d;\mu)=x(d;\mu)+iy_{\pm}(d;\mu)$,
@@
Let us focus on the case of two complex conjugate roots. The asymptotic behavior of the polynomials $\widetilde{K}_{m}(j/n)$ is given by the following lemma.
Lemma 3.
In the double scaling regime defined by
$n=HK,\quad m=gK,\quad j=\frac{K}{2}(1-2\mu),\quad K\to\infty,$ (3.5)
[…]
$(\mu-1/2)^{2}\leq g(1-g),$ (3.6)
the asymptotic of the polynomials $\widetilde{K}_{m}(i/n)$ is defined by
$\widetilde{K}_{m}(j/n)=\left(\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m))+O(1/K)\right)\cdot$
$\cdot\frac{m!}{n^{m}}\sqrt{\frac{2}{\pi K}}\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\frac{(1/2)^{m-K/2}}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}}.$ (3.7)
The functions $\theta_{\text{SDP}}^{-}$ and $\hat{\delta}(j;m)$ in the expression above is defined by
$\theta_{\text{SDP}}^{-}=-\frac{\arg(f^{\prime\prime}(t^{-}_{0}))}{2}+\frac{\pi}{2},\quad\hat{\delta}(j;m)=K\operatorname{Im}(f(t_{0}^{+}))+\arg G(t_{0}^{+}),$ (3.8)
where
$G(t)=\frac{1}{t(1-t)},\quad f(t)=(1-\mu)\ln(t-1/2)+(g-1)\ln t-g\ln(1-t),$ (3.9)
$t_{0}^{+}=\frac{1}{2\mu}\bigg(\sqrt{(g+\mu-1/2)^{2}-2g\mu}-((g-\mu)-1/2)\bigg).$ (3.10)
Proof.
For proving the lemma we will use the Proposition 4 for the integral (3.2). One can express all main values from the equation (3.3)
@@
$G(t_{0}^{\pm})=\frac{1}{t_{0}^{\pm}(1-t_{0}^{\pm})}=\frac{2\mu}{g}\sqrt{\frac{g}{1-g}}e^{\mp i\big[\zeta_{2}+2\tau\big]},$ (3.13)
@@
$\quad+\pi\cdot\Theta\left(\frac{1}{2}-\frac{\sqrt{\mu}}{2}-g\right)-\pi\cdot\Theta\left(g-\left(\frac{1}{2}+\frac{\sqrt{\mu}}{2}\right)\right)\bigg],$ (3.14)
@@
Functions $\zeta_{1}$ and $\zeta_{2}$ are the angles of these shifts relative to $\chi$,
@@
The $\sigma$ is defined from the denominator of $\pi\leq\arg f^{\prime\prime}(t_{0}^{\pm})\leq\pi$ as
@@
for $t\in\mathbb{C},\,\mu,g\in\mathbb{R},\,\mu\in(0,1),\,g\in(1/2-\sqrt{\mu(1-\mu)},1/2+\sqrt{\mu(1-\mu}))$, and
@@
The solutions of these equations are symmetric under complex conjugation $t\mapsto\overline{t}$, as a consequence of the logarithmic behavior of the action $f(t)$; their structure is illustrated in Figure 1.
Figure 1: Steepest-descent paths in the complex plane for representative values of the parameters $\mu$ and $g$. Filled dots indicate the stationary (saddle) points $t_{0}^{\pm}$ of the action $f(t)$. Arrows show the orientation of contour traversal used in the integrals. Rectangles mark two horizontal contour segments, that should be intersecting the real axis ($\operatorname{Im}t=0$) at the point $x$.
By the symmetry of the SDP branches
@@
$I=(-1)^{m}\frac{(1-p)^{m-j-K/2}}{2\pi i}\int_{\Gamma^{\prime}}g(t)e^{Kf(t)}dt\bigg|_{K\sim\infty}=$
$=(-1)^{m}(1/2)^{m-K/2}\frac{(\sqrt{1-g})^{m-1/2}}{(\sqrt{g})^{m+1/2}}\left(\sqrt{\frac{\mu}{1-\mu}}\right)^{K/2-j}\left(\sqrt{\frac{1-\mu}{1-g}}\right)^{K}\cdot$
$\cdot\sqrt{\frac{2}{\pi K}}\left(\frac{(1-\mu)\mu}{(1-g)g-\left(\mu-\frac{1}{2}\right)^{2}}\right)^{1/4}\left(\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m))+O(1/K)\right),$
@@
$I=(-1)^{m}\frac{(1-p)^{m-j-K/2}}{2\pi i}\int_{\Gamma^{\prime}}G(t)e^{Kf(t)}dt\bigg|_{K\sim\infty}=\left(\sin(\theta_{\text{SDP}}^{-}-\hat{\delta}(j;m))+O(1/K)\right)\cdot$
$=\frac{(\sqrt{1-g})^{m-K-1/2}(\sqrt{1-\mu})^{K/2+j+1/2}}{(\sqrt{g})^{m+1/2}(\sqrt{\mu})^{j-K/2-1/2}}\sqrt{\frac{2}{\pi K}}\frac{(-1)^{m}(1/2)^{m-K/2}}{\left((1-g)g-\left(\mu-\frac{1}{2}\right)^{2}\right)^{1/4}}.$
Substituting this expression to (3.1), we end the proof of the lemma. ∎
Corollary 3.
Using the equation (2.9) we can find the asymptotic behavior for $G_{m}(i/n)$ in the double scaling regime $m=gK$, $n=HK,\;K\to\infty$ for even $m$,
@@
$\cdot\bigg(\sin\big(\theta_{SDP}^{-}-\hat{\delta}(i;m+2)\big)+\frac{(K-m)}{m+2}\frac{g}{1-g}\sin\big(\theta_{SDP}^{-}-\hat{\delta}(i;m)\big)+O(1/K)\bigg).$ (3.18)
4 Limiting behavior of the correlation kernel
The main goal of this section is to give a proof of the Theorem 1. To do that, we compute the asymptotic of the correlation kernel (1.14) in following regime of the symplectic polynomials
@@ 4 Limiting behavior of the correlation kernel
with $K\to\infty$ and fixed $g,H$, and $m$ is the number of polynomial.
Proof of Theorem 1..
The equation (3.18) for $m=2n$ has the following form
$G_{2n}(i/n)=g_{2n}(i/n)\sqrt{\Lambda_{2n}}=\frac{1}{(i/n)^{2}}\frac{(2n+2)!}{n^{2n+2}}F_{2n}(i)\cdot$
$\bigg(\sin\big(\theta_{SDP}^{-}-\hat{\delta}(i;2n+2)\big)+\sin\big(\theta_{SDP}^{-}-\hat{\delta}(i;2n)\big)+O(1/K)\bigg),$ (4.2)
@@
with $g=2H$ and in general $0<H<1/2$ with $H=1/4$ for $n=k$ (square Young tableaux). Now we substitute asymptotic of the polynomials into the kernel to demonstrate its convergence to the discrete sine-kernel
@@
$\cdot\bigg(\sin\big(\theta^{-}_{\text{SDP}}(j;2n)-\hat{\delta}(j;2n)\big)+\sin\big(\theta^{-}_{\text{SDP}}(j;2n-2)-\hat{\delta}(j;2n-2)\big)\bigg)-[ij]\bigg)+O(1/K)\Bigg),$ (4.3)
@@
we aim to derive the asymptotic of $\lambda$ in the bulk regime for 1/2<x<1/2-1/2<x<1/2
@@
$\gamma,\hat{\zeta}_{1},\hat{\zeta}_{2},\hat{\tau},\arg f^{\prime\prime}(t_{0}^{-}))$. We need to choose the interval between the critical points of the action. So we can expand the piecewise functions for example in $g>1/2,\,\mu>1/2$ (K/2<i<0-K/2<i<0)
@@
$\arg f^{\prime\prime}(t_{0}^{-})=\tan^{-1}\left(\frac{(2g-1)(2\mu-1)}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+$
$+2\tan^{-1}\left(\frac{-\frac{\mu}{g-1/2}+2g-1}{\sqrt{-4(g-1)g-(1-2\mu)^{2}}}\right)+\pi.$
In the other domains of parameter space such as $g>1/2,\mu<1/2$ we will get the same formulas up to $\pm\pi$. These contributions will cancel each other out in final answer. Next we derive the asymptotic of trigonometric part of the kernel accurate up to $O(1/K)$. It is obvious that the shift $m\to m+2$ generate the shift $g\to g+2/K$, so, by the definition of $\lambda(i;2n)$
@@
$\beta=\left(-2\zeta_{2}-\frac{\partial\hat{\zeta}_{1}}{\partial g}2+\frac{\partial\gamma}{\partial g}\right)=2\cot^{-1}\left(\frac{x}{\sqrt{2H-4H^{2}-x^{2}}}\right)+O(1/K),$
@@
$=\bigg(\sin\gamma_{i}+\sin\gamma_{i}\cos\beta+\cos\gamma_{i}\sin\beta\bigg)\bigg(\sin\gamma_{j}+\sin\gamma_{j}\cos\beta-\cos\gamma_{j}\sin\beta\bigg)$-
$-[ij]+O(1/K)=-2\sin\beta(1+\cos\beta)\sin(\gamma_{i}-\gamma_{j}),$
@@
The normalization of the kernel is not important because it can be cancelled out by the renormalization of the polynomials. Therefore we have established the pointwise convergence of the Christoffel–Darboux kernel to the discrete sine-kernel
@@
which ends the proof of Theorem 1. ∎
The limit density coincides with the result for the limit shape of symplectic diagrams obtained in [nazarov2024skew, Thm. 5.1] after substituting $c=\lim\frac{k}{n}=\frac{1-2H}{2H}$.
@@
Figure 2: Plot of the ratio $\frac{\mathcal{K}(i,j)}{\mathcal{K}(i,i)}$ for $i=75$ obtained by sampling 500500 samples, each of $10000$ random Young diagrams inside of $n\times k$ box with $n=50,k=100$ using Proctor’s algorithm (box and whiskers plot) and its comparison to the same ratio for the Christoffel-Darboux kernel (1.10) (dashed black line) and discrete sine kernel (dotted blue) (4.4), (4.5)
5 Conclusion and future work